26' A vehicle starting its motion from rest travels at a uniform acceleration of 5ms-2 for a certain time and



travels with the attained uniform velocity for another certain period of time. Then it decelerates uniformly at



5ms-2 and stops after a certain time. If the total time spent is 25s and the average velocity for the whole



motion is 20ms-1 , find the time that the vehicle travelled at uniform velocity.

1
Expert's answer
2022-02-06T14:33:29-0500

Given:

"v_0=0"

"a_1=\rm 5\: m/s^2"

"t=25\:\rm s"

"v_{\rm ave}=20\:\rm m/s"


Let us denote "\tau" the time that the vehicle travelled at uniform velocity. Then, the total distance rtaveled by vehicle

"d=a((25-\tau)/2)^2+a(25-\tau)/2\tau"

The average velocity

"v_{\rm ave}=\frac{d}{t}=\frac{a((25-\tau)/2)^2+a(25-\tau)/2\tau}{t}"

We get an equation

"\frac{5((25-\tau)/2)^2+5(25-\tau)/2\tau}{25}=20"

"((25-\tau)/2)^2+\tau(25-\tau)/2=100"

"(25-\tau)^2+2\tau(25-\tau)=400"

"\tau^2=75"

"\tau=5\sqrt{3}\:\rm s"